# Federated Semi-Supervised and Semi-Asynchronous Learning for Anomaly Detection in IoT Networks

### v2

@@ IV-C1 Semi-Asynchronous Model Update
First, we design a semi-asynchronous model update approach that utilizes a hyperparameter $C$ to control the proportion of clients participating in each round of the global aggregation. This idea has been widely applied to anomaly detection in IoT networks, for instance, the most existing methods [8, 9, 10, 11] use FedAvg [7] for the global model update, whose idea is to select the $C$ proportion of clients in advance to participate in each round of training, and then wait for all selected clients to complete the training before the global aggregation, which is essentially a synchronous method.
@@ IV-C2 Staleness Tolerant Distribution
With the combination of the semi-asynchronous model update and staleness tolerant distribution, we achieve the version control of the local models. We synchronize the latest clients with the server to prevent model divergence. At the same time, deprecated clients are forced to update, so that the global model will not be poisoned by the heavily outdated local models. Furthermore, we allow tolerant clients to remain asynchronous with the server, so that the data and knowledge of each client can be fully utilized to improve the performance. Model staleness control has been shown to be the key to guaranteeing convergence [32]. Poorly controlled staleness, such as being too small, can cause some slower clients to have large staleness before completing their local training. They will be easily forced to synchronize. Therefore, these clients may never participate in global updates, which reduces the training accuracy, especially on non-IID data.
@@
For instance, as shown in Fig. 3, there are five clients $C_{1}\sim C_{5}$ and a server $S$ with $C=0.4,\tau=2$. Clients perform unsupervised learning, while the server performs supervised learning. At the round $r_{0}$, clients $C_{1}$ and $C_{2}$ are the first to complete local training and upload model parameters $\omega_{1}^{1}$ and $\omega_{2}^{1}$ to the server $S$. Since the server has received the parameters from $C*M=0.4*5=2$ clients, the new global model $\omega_{g}^{1}$ is updated by these two unsupervised parameters and the supervised parameter of the server $\omega_{s}^{1}$. Meanwhile, the staleness of each client at different rounds is shown in Table II. The staleness of $C_{1}$ and $C_{2}$ is 00 due to the participation at this round while the staleness of others is $21<\tau=$ 2, which means they are tolerable clients and do not need to be forced to update. Then, at the round $r_{1}$, $C_{1}$ and $C_{3}$ participate in the global update, so their staleness is updated to 00, while the staleness of other clients is increased by one. Next, at the round $r_{2}$, $C_{2}$ and $C_{4}$ successfully complete the local training. Meanwhile, it is worth noting that the staleness of $C_{5}$ becomes $3>2$, which may be due to the network failure or system crash. Therefore, $C_{5}$ is a deprecated client and the server needs to distribute the latest global model $\omega_{g}^{3}$ to it. Finally, $C_{5}$ successfully participates in the global update at the round $r_{3}$.
@@ IV-D Design of Aggregation Function
Most of the existing FL-based anomaly detection schemes for IoT networks use FedAvg as the aggregation function, which is the most dominant method and has been proven to be convergent [41]. In this paper, since we consider a novel semi-supervised learning and semi-asynchronous model update scenario, which is different from that of the existing works, a new aggregation function should be proposed. A naive approach is to simply transfer FedAvg over, as shown below.
@@ IV-D3 Group-based Aggregation Function
The non-IID nature of data leads to the model diversity among different clients, which brings significant difficulties to the training and aggregation [42]. To address this issue, we propose a group-based aggregation function for the robust optimization and convergence of the model. We utilize the K-means clustering method to divide clients into $|G|$ groups based on the data distribution of clients, and then perform the global aggregation. Specifically, after clients complete the FSSL training and upload their local models in a semi-asynchronous manner, the server divide them into $|G|$ groups, and then updates the global model according to the following formula.
@@ IV-E Adaptive Learning Rate
where $n_{i}$ is the number of times that the client $c_{i}$ has participated in the global update. All past rounds are regarded equally and given the same weight, which is natural but suffers from an ill-consideration. Taking Fig. 3 as an example, the client $C_{1}$ participates the global update at the rounds $r_{0}$ and $r_{1}$, $C_{2}$ participates the global update at the rounds $r_{0}$ and $r_{2}$, and $C_{3}$ participates the global update at the rounds $r_{1}$ and $r_{3}$. All three of them participate in two rounds of global updates, whose frequencies will be considered the same according to the above formula. However, intuitively, the influence of the local model $\omega_{i}^{r^{\prime}+1}$ uploaded by the client $C_{i}$ at the round $r^{\prime}$ on the global model $\omega_{g}^{r+1}$ gradually decays with the increase of the current training round $r$. In other words, the global model will gradually forget knowledge from local models uploaded in the previous rounds and be replaced by the knowledge in more recent rounds. As a result, the more recent the participation rounds of the client and the newer the historical rounds of the client, the greater the impact on the global model. The impacts of $C_{1}$, $C_{2}$ and $C_{3}$ on the global model, in descending order, should be as follows: $C_{3},C_{2},C_{1}$. Therefore, more recent rounds need to be given greater weights while older rounds will be given lower weights. There are many round-weight functions which can satisfy this requirement, such as $h(r)=(1+a)^{r},a>0$, the one inspired by the exponential smoothing [43].
@@ IV-F Communication Cost Optimization
In federated learning for anomaly detection in IoT networks, the communication cost is mainly related to two parameters, one is the data size of exchanged model parameters at each round, and the other is the number of rounds required for model convergence. On the one hand, the communication cost can be reduced by increasing the number of epochs for each local training. However, as the number of epochs increases, the diversity between different local models also increase, especially on non-IID data. On the other hand, model compression and reconstruction techniques such as structured updates and sketched updates [44] are exploited to reduce the exchanged data size per round. However, many compression methods also inevitably exacerbate the error of local model parameters and the diversity between different models. The aggregation of these diverse models may result in a decrease in the accuracy of model performance and speed of model convergence.
@@ V-A Dataset and Scenarios
In this paper, we use CIC-IDS 2017 [22] dataset for experimental evaluation, as it is one of the largest datasets in the anomaly detection field. It can well reflect the real network environment, and has been widely used. The dataset contains 78-dimension attribute features and 1-dimension classification feature which contains 15 attack types. We discard several attacks with very sparse data, and the final experimental dataset contains 8 attack types: DoS Hulk, PortScan, DDoS, DoS GoldenEye, FTP-Patator, SSH-Patator, DoS slowloris, and DoS Slowhttp. The whole dataset has about 3 million data, however, due to the extremely unbalanced distribution of categorizes, that is about 75 percentage of the data is benign, we select about 540,000 data for experiment, of which the ratio of training and test-validation data is about 9:1. Meanwhile, the labeled training data of the server accounts for about 5 percentage of the total training data.
Furthermore, in order to evaluate the performance of our scheme under different data distributions, we design two different scenarios, namely basic and balanced scenarios, whose data distributions are shown in Table III. In the basic scenario, the distribution of classes and samples among different clients is highly unbalanced. It represents a real IoT network in which some devices work normally and perform the expected operations while others suffer from multiple attacks. In the balanced scenario, the data of each client is independent and identically distributed, and only the amount of their data is different. Meanwhile, Shannon Entropy [45] is used to measure the imbalance of the local data of each client $C_{i}$, which can be given by
@@ V-B Experiment Setup
We use two mainstream machine learning frameworks, Keras Library and TensorFlow backend [46] to implement our experiments, that employs a federated architecture for learning. The system is set up with a security service provider (i.e., the server) and 10 security gateways (i.e., clients), each of which deploys an anomaly detection component and monitors several heterogeneous IoT devices. The CNN model is used for anomaly detection which consists of two 1D-CNN (the filters are 128 and 256 respectively), one flatten, one fully-connected (the number of hidden neurons is 256 and the activation function is ReLU), one dropout (the dropout rate is 0.1) and one fully-connected (the activation function is Softmax) layers. Table IV summarizes other default experimental settings.
@@ V-F1 Benchmarks
    FedAvg-SSL: FedAvg [7] is a mainstream federated learning method and is widely used for anomaly detection in IoT networks [8, 9, 10, 11]. Based on the participation proportion of clients, it can be divided into two types, FedAvg-SSL-Partial and FedAvg-SSL-All. In FedAvg-SSL-Partial, at each round, the server randomly pre-specify partial clients to participate in this round of FL training process, while in FedAvg-SSL-All, all clients should participate in each round of training. In this paper, for consistency, we set the number of participating clients per round for FedAvg-SSL-Partial to 6.
@@
    FedAsync-SSL: FedAsync is introduced in [23] which is a classic asynchronous federated learning algorithm, and its idea has been applied to anomaly detection in IoT networks [12]. The server starts the global update once it receives one local model from an arbitrary client. In this paper, we use the polynomial function as the staleness function and set $\alpha=0.9$, $a=0.5$, $t-\tau\leq 16$ and $\rho=0.005$, which is claimed to be the combination with the best performance in [23].
@@
    Local-SSL: At the same time, Local-SSL [26] aggregates the labeled data of the server and the unlabeled data of all clients to perform the local semi-supervised training, which should be the ceiling of the performance. It is used to evaluate and measure the gap between our distributed FedS3A method and the centralized Local-SSL.

### v3

@@ IV-C1 Semi-Asynchronous Model Update
First, we design a semi-asynchronous model update approach that utilizes a hyperparameter $C$ to control the proportion of clients participating in each round of the global aggregation. This idea has been widely applied to anomaly detection in IoT networks, for instance, the most existing methods [DIoT, NBaIoTCN, ToNIoTCN, DADIoT] use FedAvg [FedAvg] for the global model update, whose idea is to select the $C$ proportion of clients in advance to participate in each round of training, and then wait for all selected clients to complete the training before the global aggregation, which is essentially a synchronous method.
@@ IV-C2 Staleness Tolerant Distribution
With the combination of the semi-asynchronous model update and staleness tolerant distribution, we achieve the version control of the local models. We synchronize the latest clients with the server to prevent model divergence. At the same time, deprecated clients are forced to update, so that the global model will not be poisoned by the heavily outdated local models. Furthermore, we allow tolerant clients to remain asynchronous with the server, so that the data and knowledge of each client can be fully utilized to improve the performance. Model staleness control has been shown to be the key to guaranteeing convergence [FedAT]. Poorly controlled staleness, such as being too small, can cause some slower clients to have large staleness before completing their local training. They will be easily forced to synchronize. Therefore, these clients may never participate in global updates, which reduces the training accuracy, especially on non-IID data.
@@
For instance, as shown in Fig. 3, there are five clients $C_{1}\sim C_{5}$ and a server $S$ with $C=0.4,\tau=2$. Clients perform unsupervised learning, while the server performs supervised learning. At the round $r_{0}$, clients $C_{1}$ and $C_{2}$ are the first to complete local training and upload model parameters $\omega_{1}^{1}$ and $\omega_{2}^{1}$ to the server $S$. Since the server has received the parameters from CM=0.45=2C*M=0.4*5=2 clients, the new global model $\omega_{g}^{1}$ is updated by these two unsupervised parameters and the supervised parameter of the server $\omega_{s}^{1}$. Meanwhile, the staleness of each client at different rounds is shown in Table II. The staleness of $C_{1}$ and $C_{2}$ is 0 due to the participation at this round while the staleness of others is $1<\tau=2$, which means they are tolerable clients and do not need to be forced to update. Then, at the round $r_{1}$, $C_{1}$ and $C_{3}$ participate in the global update, so their staleness is updated to 0, while the staleness of other clients is increased by one. Next, at the round $r_{2}$, $C_{2}$ and $C_{4}$ successfully complete the local training. Meanwhile, it is worth noting that the staleness of $C_{5}$ becomes $3>2$, which may be due to the network failure or system crash. Therefore, $C_{5}$ is a deprecated client and the server needs to distribute the latest global model $\omega_{g}^{3}$ to it. Finally, $C_{5}$ successfully participates in the global update at the round $r_{3}$.
@@ IV-D Design of Aggregation Function
Most of the existing FL-based anomaly detection schemes for IoT networks use FedAvg as the aggregation function, which is the most dominant method and has been proven to be convergent [FedAvgConPro]. In this paper, since we consider a novel semi-supervised learning and semi-asynchronous model update scenario, which is different from that of the existing works, a new aggregation function should be proposed. A naive approach is to simply transfer FedAvg over, as shown below.
@@ IV-D3 Group-based Aggregation Function
The non-IID nature of data leads to the model diversity among different clients, which brings significant difficulties to the training and aggregation [Group-Based]. To address this issue, we propose a group-based aggregation function for the robust optimization and convergence of the model. We utilize the K-means clustering method to divide clients into $|G|$ groups based on the data distribution of clients, and then perform the global aggregation. Specifically, after clients complete the FSSL training and upload their local models in a semi-asynchronous manner, the server divide them into $|G|$ groups, and then updates the global model according to the following formula.
@@ IV-E Adaptive Learning Rate
where $n_{i}$ is the number of times that the client $c_{i}$ has participated in the global update. All past rounds are regarded equally and given the same weight, which is natural but suffers from an ill-consideration. Taking Fig. 3 as an example, the client $C_{1}$ participates the global update at the rounds $r_{0}$ and $r_{1}$, $C_{2}$ participates the global update at the rounds $r_{0}$ and $r_{2}$, and $C_{3}$ participates the global update at the rounds $r_{1}$ and $r_{3}$. All three of them participate in two rounds of global updates, whose frequencies will be considered the same according to the above formula. However, intuitively, the influence of the local model $\omega_{i}^{r^{\prime}+1}$ uploaded by the client $C_{i}$ at the round $r^{\prime}$ on the global model $\omega_{g}^{r+1}$ gradually decays with the increase of the current training round $r$. In other words, the global model will gradually forget knowledge from local models uploaded in the previous rounds and be replaced by the knowledge in more recent rounds. As a result, the more recent the participation rounds of the client and the newer the historical rounds of the client, the greater the impact on the global model. The impacts of $C_{1}$, $C_{2}$ and $C_{3}$ on the global model, in descending order, should be as follows: $C_{3},C_{2},C_{1}$. Therefore, more recent rounds need to be given greater weights while older rounds will be given lower weights. There are many round-weight functions which can satisfy this requirement, such as $h(r)=(1+a)^{r},a>0$, the one inspired by the exponential smoothing [ES].
@@ IV-F Communication Cost Optimization
In federated learning for anomaly detection in IoT networks, the communication cost is mainly related to two parameters, one is the data size of exchanged model parameters at each round, and the other is the number of rounds required for model convergence. On the one hand, the communication cost can be reduced by increasing the number of epochs for each local training. However, as the number of epochs increases, the diversity between different local models also increase, especially on non-IID data. On the other hand, model compression and reconstruction techniques such as structured updates and sketched updates [Communication] are exploited to reduce the exchanged data size per round. However, many compression methods also inevitably exacerbate the error of local model parameters and the diversity between different models. The aggregation of these diverse models may result in a decrease in the accuracy of model performance and speed of model convergence.
@@ V-A Dataset and Scenarios
In this paper, we use CIC-IDS 2017 [CIC-IDS] dataset for experimental evaluation, as it is one of the largest datasets in the anomaly detection field. It can well reflect the real network environment, and has been widely used. The dataset contains 78-dimension attribute features and 1-dimension classification feature which contains 15 attack types. We discard several attacks with very sparse data, and the final experimental dataset contains 8 attack types: DoS Hulk, PortScan, DDoS, DoS GoldenEye, FTP-Patator, SSH-Patator, DoS slowloris, and DoS Slowhttp. The whole dataset has about 3 million data, however, due to the extremely unbalanced distribution of categorizes, that is about 75 percentage of the data is benign, we select about 540,000 data for experiment, of which the ratio of training and test-validation data is about 9:1. Meanwhile, the labeled training data of the server accounts for about 5 percentage of the total training data.
Furthermore, in order to evaluate the performance of our scheme under different data distributions, we design two different scenarios, namely basic and balanced scenarios, whose data distributions are shown in Table III. In the basic scenario, the distribution of classes and samples among different clients is highly unbalanced. It represents a real IoT network in which some devices work normally and perform the expected operations while others suffer from multiple attacks. In the balanced scenario, the data of each client is independent and identically distributed, and only the amount of their data is different. Meanwhile, Shannon Entropy [Shannon] is used to measure the imbalance of the local data of each client $C_{i}$, which can be given by
@@ V-B Experiment Setup
We use two mainstream machine learning frameworks, Keras Library and TensorFlow backend [Keras] to implement our experiments, that employs a federated architecture for learning. The system is set up with a security service provider (i.e., the server) and 10 security gateways (i.e., clients), each of which deploys an anomaly detection component and monitors several heterogeneous IoT devices. The CNN model is used for anomaly detection which consists of two 1D-CNN (the filters are 128 and 256 respectively), one flatten, one fully-connected (the number of hidden neurons is 256 and the activation function is ReLU), one dropout (the dropout rate is 0.1) and one fully-connected (the activation function is Softmax) layers. Table IV summarizes other default experimental settings.
@@ V-F1 Benchmarks
    FedAvg-SSL: FedAvg [FedAvg] is a mainstream federated learning method and is widely used for anomaly detection in IoT networks [DIoT, NBaIoTCN, ToNIoTCN, DADIoT]. Based on the participation proportion of clients, it can be divided into two types, FedAvg-SSL-Partial and FedAvg-SSL-All. In FedAvg-SSL-Partial, at each round, the server randomly pre-specify partial clients to participate in this round of FL training process, while in FedAvg-SSL-All, all clients should participate in each round of training. In this paper, for consistency, we set the number of participating clients per round for FedAvg-SSL-Partial to 6.
@@
    FedAsync-SSL: FedAsync is introduced in [AFL] which is a classic asynchronous federated learning algorithm, and its idea has been applied to anomaly detection in IoT networks [AFLIoT]. The server starts the global update once it receives one local model from an arbitrary client. In this paper, we use the polynomial function as the staleness function and set $\alpha=0.9$, $a=0.5$, $t-\tau\leq 16$ and $\rho=0.005$, which is claimed to be the combination with the best performance in [AFL].
@@
    Local-SSL: At the same time, Local-SSL [Fixmatch] aggregates the labeled data of the server and the unlabeled data of all clients to perform the local semi-supervised training, which should be the ceiling of the performance. It is used to evaluate and measure the gap between our distributed FedS3A method and the centralized Local-SSL.
@@ Acknowledgment
[Uncaptioned image] Wenbin Zhai is currently pursing the Ph.D. degree with the Department of Computing, The Hong Kong Polytechnic University. He received the B.S. degree from Nanjing University of Chinese Medicine, Nanjing, Jiangsu Province, China in 2020, and the M.S. degree in Nanjing University of Aeronautics and Astronautics, Nanjing, Jiangsu Province, China in 2023. His research interests include AI security, Cybersecurity, Wireless Sensor Networks (WSNs), and Networking.
[Uncaptioned image] Liang Liu is currently an associate professor in College of Computer Science and Technology, Nanjing University of Aeronautics and Astronautics, Nanjing, Jiangsu Province, China. His research interests include distributed system, big data and system security. He received the B.S. degree in computer science from Northwestern Polytechnical University, Xi’an, Shanxi Province, China in 2005, and the Ph.D. degree in computer science from Nanjing University of Aeronautics and Astronautics, Nanjing, Jiangsu Province, China in 2012.
[Uncaptioned image] Feng Wang graduated from Zhicheng College of Fuzhou University with a bachelor’s degree in 2018. At present, he is studying for a master’s degree in the collage of Computer Science and Technology, Nanjing University of Aeronautics and Astronautics. His main research direction is moving target defense and distributed databases.
[Uncaptioned image] Youwei Ding is currently a lecturer in the School of Artificial Intelligence and Information Technology, Nanjing University of Chinese Medicine, Nanjing, Jiangsu Province, China. His research interests include energy efficient data management, big data analysis and data security. He received the B.S and M.S degrees in computer science from Yangzhou University, Yangzhou, Jiangsu Province, China in 2007 and 2010, and the Ph.D. degree in computer science form Nanjing University of Aeronautics and Astronautics, Nanjing, Jiangsu Province, China in 2016.
[Uncaptioned image] Wanying Lu graduated from Henan Polytechnic University with a bachelor’s degree in 2021. At present, she is studying for a master’s degree in the collage of Computer Science and Technology, Nanjing University of Aeronautics and Astronautics. Her main research direction is time series big data storage and data mining.
[Uncaptioned image] Weizhi Meng received the Ph.D. degree in computer science from the City University of Hong Kong, Hong Kong, in 2013. He is currently a Professor with the School of Computing and Communications, Lancaster University, Lancaster, U.K. He was a Faculty Member with the Department of Applied Mathematics and Computer Science, Technical University of Denmark, Denmark. His primary research interests include cyber security and intelligent technology in security, including intrusion detection, smartphone security, biometric authentication, HCI security, cloud security, trust management, blockchain in security, cyber–physical system security, and IoT security. Prof. Meng won the Outstanding Academic Performance Award during his doctoral study, the IEEE ComSoc Best Young Researcher Award for Europe, Middle East, and Africa Region in 2020, and the IEEE MGA Young Professionals Achievement Award in 2020. He is a recipient of the Hong Kong Institution of Engineers Outstanding Paper Award for Young Engineers/Researchers in 2014 and 2017.